# IIb or not IIb: A Catalog of ZTF Kilonova Imposters

### v1

@@ 1 Introduction
KNe, which are powered by the thermal heating caused by the radioactive decay of newly synthesized and unstable r-elements (Metzger, 2019), can provide smoking gun evidence of neutron star mergers. This thermal emission peaks around a day after the initial merger, followed by a rapid fade on the order of less than a week (Metzger et al., 2010). In addition to the aforementioned electromagnetic (EM) signal, these mergers can also produce a $\gamma$-ray burst (GRB) and are detectable via gravitational waves (GW) generated from the compact binary coalescence (CBC) (Metzger et al., 2010; Jin et al., 2016; Lehner et al., 2016; Tanaka, 2016). It was through detections by the International Gravitational Wave Network (IGWN) of the binary neutron star merger GW170817 (Abbott et al., 2017a) that the first (and only) KN associated with a GW signal, AT2017gfo, was localized and subsequently observed in a spate of multi-wavelength observations (Coulter et al., 2017; Smartt et al., 2017; Abbott et al., 2017a) in addition to the associated GRB170817A (Goldstein et al., 2017; Savchenko et al., 2017; Abbott et al., 2017b).
@@ 4 Data Processing
In the aggregate, all of the transient types in this study are characterized by a rise followed by a fall. Due to the nature of transient observation cadences, a smaller number of objects were observed during their rise compared to the number observed during their fade. There are 3 objects observed only during their rise, 36 objects observed only during their fade, and 30 objects observed for both their rise and fade. The distribution of object types in the sample is shown in Figure 2. A brief description of each of these classes can be found in Appendix A.
@@
Observational data is cut to center around the rise, peak, and fade. Of the 33 objects with an observed rise, 8 have a rise duration that exceeds 10 days in any filter; of these, 5 are SNe IIb, and 3 are SNe Ia. For the purposes of this paper, we measure the fade as the difference in magnitude between the peak luminosity and observations over the next 2 days. This is because KNe, as fast fading objects, would generally only be visible for a few days following their peak. An object that is still observable by ZTF for weeks following its peak is unlikely to be a KN.
@@ 4.1 Gaussian Process
The GP is applied to the first 50 days of data for all objects in our sample [^2] We restrict our analysis to the first 50 days because this study largely focuses on the rapid rise and fade of transients, making longer term data unnecessary.. This results in smooth light curves with high observing cadences, allowing for more precise analysis of rapid changes in luminosity. From there, each of the objects was reduced to only include observations at and around the peak luminosity, and the flux values from the GP are converted back to magnitude space by inverting Equations 1 and 2. An example of the GP applied to one of the objects is shown in Figure 10.
After applying the GP to the 73 objects with classifications and redshifts, 4 of the transients did not have sufficient data to be modeled by the GP; an example of one such object can be seen in Figure 9, which the GP models as essentially flat. As a result, the total number of objects included in the study is 69, with 11 different transient types represented in the sample. The distribution of types in the sample is shown in Figure 2. A full list of objects can be found in Table LABEL:tab:objects.
@@ 5 Evaluation Metrics
We compare the sample of transients to a grid of kilonova models computed with the 3D Monte Carlo radiative transfer code POSSIS (Bulla, 2019, 2023). In particular, we use the grid presented in (Dietrich et al., 2020b) [^3] The grid is publicly available at https://bit.ly/possis_models. The ejecta are divided into two main components: a disk-wind ejecta component in the inner region (i.e., low velocities) and a dynamical ejecta component in the outer regions (i.e., high velocities) characterized by lanthanide-rich compositions close to the merger plane and lanthanide-poor compositions at polar angles. The masses of the two components, $m_{\rm wind}$ and $m_{\rm dyn}$, and the half-opening angle of the lanthanide-rich component, $\phi$, are varied to create a grid of 196 models, with observables extracted for $11$ different viewing angles. The values for the model grid are generally plotted as black points in all figures.
We also calculate these metrics for AT2017gfo and include it under the “AT2017gfo” label. We note that, since ZTF was not operating at the time of its detection, there are no ZTF observations of AT2017gfo. Instead, we use observations from Pan-STARRS, which captured the fade in 3 bands similar to those of ZTF. The same GP is applied to these observations, which can be seen in Figure 8. Across all the metrics evaluated, AT2017gfo generally falls within the same range as the KN model grid.
In addition to the following sections, which consider a combination of metrics, we evaluate the numerical separation between transient classes and the model grid with respect to each metric using a ”distance” method that is described in Section C.
@@ 5.1 Rise and Fade Rates
Figure 4: Fade rate vs/ peak luminosity in each filter for the transient sample compared to the KNe model grid. Generally more objects were observed during their fade as compared to their rise (See Figure 3). Most of the transients are clustered just above the KNe model grid with comparable fade rates. The model grid demonstrates two populations distinguished by fade rate, with a gap between the faster and slower fading populations that is most noticeable in the g-band.
@@ 6 Conclusions
While there is variance in the qualitative performance of each metric, most of the metrics discussed in this paper show a mild discriminatory utility; using these metrics in tandem with some sort of weighting could be used to calculate a second order metric to indicate the likelihood of a transient being a KN candidate. This would return results on a timescale shorter than inference-based methods that use tools like NMMA (Barna et al., 2024).
@@ A.1 GRB Afterglows
GRB afterglows (referred to as afterglows for much of this paper) occur as a result of the shock originating from a GRB interacting with the interstellar medium in the form of synchrotron emission (Sari et al., 1998).
Like KNe, afterglows decline rapidly (Rhoads, 1999), and afterglows occur from a similar mechanism to the models proposed for KNe; indeed, CBCs give rise to both a KN and a short GRB (sGRB) afterglow, as was observed with GW170817 and its counterparts (Abbott et al., 2017b). In these cases, it can be difficult to disentangle the two signals from each other given the greater brightness of afterglows (Zhu et al., 2022), though multi-wavelength observations can mitigate this challenge (Wallace & Sarin, 2024).
@@ A.5 Tidal Disruption Events
Tidal disruption events (TDEs) occur when a star orbiting a supermassive black hole (SMBH) is drawn so close to the SMBH that it is pulled apart by tidal forces. This phenomenon was first theorized in relation to quasi-stellar sources (QSOs) (Hills, 1975). While the first observations of TDEs occurred in the x-ray regime (Piro et al., 1988; Donley et al., 2002), TDEs are observable across the EM spectrum depending on the viewing angle (Dai et al., 2018). A number of TDEs have been detected by ZTF (van Velzen et al., 2021), making them a potential contaminant, but their presence could be mitigated with galaxy correlation given their presence at the center of galactic nuclei and their higher peak luminosity in comparison to KNe (Lodato & Rossi, 2011)
A.6 SNe Ia
Type Ia supernovae (SNe Ia) are the only class of SN that undergoes thermonuclear runaway for its explosion mechanism. This occurs when a WD in a binary system accretes enough mass to hit the Chandrasekhar limit, which leads to the thermonuclear runaway (Chandrasekhar, 1931; Baym et al., 1971). Due to this roughly similar explosion mass, SNe Ia are understood to have standardizable peak luminosities (Phillips, 1993), though there are a subset of SNe Ia that are under-luminous, which are referred to as SNe Iax (Foley et al., 2013).
@@ A.6 SNe Ia
A.7 SNe Ib
Like SNe Ia, SNe Ib are categorized as Type I SNe due to the paucity of hydrogen in their spectra, but SNe Ib are driven by core collapse (CC), which occurs when the core of the star is no longer able to counter the gravitational forces of its own mass (Clocchiatti & Wheeler, 1997). CC SNe progenitors are massive stars, generally of the order 8 $M_{\odot}$ at minimum (Heger et al., 2003; Díaz-Rodríguez et al., 2021), with some debate as to the upper mass limit (Smartt, 2009, 2015; Sukhbold et al., 2016; Kuncarayakti et al., 2018; Díaz-Rodríguez et al., 2021).
@@ A.8 SNe Ibn
A.9 SNe II
Like SNe Ib, SNe II occur as a result of the CC of massive stars, but they differ in that their red supergiant progenitors retain their hydrogen envelope (Chevalier & Fransson, 1994; Filippenko, 1997). There are a number of subclasses of SNe II, of which two are described below.
@@ A.10 SNe IIP
A.11 SNe IIb
SNe IIb are a unique class of CC SNe that initially appear similar to SNe II but later evolve to resemble SNe Ib (Filippenko, 1997). This occurs due to a milder stripping of the outer hydrogen layers at early times in comparison to other SE SNe, with the associated hydrogen spectral features disappearing at later times, coming to resemble the spectra of SNe Ib (Sravan et al., 2019). A unique observational trait of SNe IIb is their double-peak feature (Crawford et al., 2025). The initial peak is powered by shock-cooling emission (Kamble et al., 2016; Armstrong et al., 2021), which later gives way to the secondary peak that is driven by the decay of 56Ni produced during the initial explosion (Arnett, 1980; Pellegrino et al., 2023). The first peak demonstrates a more rapid rise and fade than the second, but both peaks fade faster than something like a SN IIP. SNe IIb represent the largest subclass of observed SE SNe (Li et al., 2011a; Shivvers et al., 2017), though they are generally overrepresented in our sample relative to SNe IIP, which are the most observed subclass of SNe generally (Shivvers et al., 2017), suggesting they are a major contaminant in KNe searches.
Appendix B GP Interpolation
Not all of the objects with known redshifts were suited for the GP. Figure 9 shows ZTF20abwysqy, a long GRB, which was not well interpolated by the GP. This is due to the relatively low number of observations, which take place over a single night. The third data point in the g-band suggests that it might be demonstrating a fade, but the error is too large to say this definitively. Regardless, this shows that there is a lower limit in terms of observations below which the GP cannot effectively interpolate the data.
Conversely, Figure 10 shows a IIb SN, ZTF20abwysqy, that was well fit by the GP. While this is much more observed than ZTF20abwysqy, this is an unevenly sampled lightcurve with a somewhat ambiguous peak in the g-band and a paucity of observations immediately following the peak in the r-band, which would make calculating several metrics, including the color at peak and FWHM, less consistent. The GP creates a plausible interpolation of the observed data, which enables more consistent calculations of metrics like the fade rate and FWHM.
Appendix C Metric Distance
$d=\frac{\sum_{i=0}^{n_{type}}\sum_{i=0}^{n_{grid}}|x_{i}-x_{j}|}{n_{type}\times n_{grid}}$ (C1)
For visual clarity, we condense the imposters into 5 groups
  • GRB afterglows
  • IIb SNe
  • Misc SNe (Ia, Ib, Ibn, II, IIp)
  • Galactic transients (CVs, novae)
  • Other transients (FBOTs, TDEs)
Figure 12 shows radar plots for all of the metrics discussed in Section 5. Take the first row, which corresponds to the peak luminosity, as a test case. The KN model has a fairly narrow range in terms of peak luminosity, so the self similarity value is almost zero. Across all filters, galactic sources show the highest value, which is logical given that these transients are known to have much lower peak absolute luminosities in comparison to KNe and other extragalactic transients; however, they can be a contaminant due to their relative proximity resulting in a high apparent magnitude. One should also note that, while it is fairly small, AT2017gfo has a value slightly above that of the KN model baseline, which is most pronounced in the g-band.
The fourth row shows the FWHM for each filter. As shown in Figures 6 and 7, the range of FWHM values is fairly narrow for the model grid; correspondingly, the baseline metric distance value is relatively low. The transient classes demonstrating the greatest distance from the baseline are both categories of SNe, with IIb SNe being edged out by the other collected SNe category. SNe occur on a longer timescale in comparison to KNe, and this is reflected in Figure 6. The galactic and miscellaneous sources are also slightly above the baseline value, but much less so.
This metric distance approach demonstrates a plausible level of discriminatory power using basic metrics that can be quickly and easily calculated; the authors of this paper invite readers to experiment with second order metrics and evaluate their performance in relation to discriminatory power.
Appendix D Transient List

### v2

@@ 1 Introduction
Kilonovae, which are powered by the thermal heating caused by the radioactive decay of newly synthesized and unstable r-elements (Li & Paczyński, 1998; Metzger, 2019), can provide smoking gun evidence of neutron star mergers. This thermal emission peaks around a day after the initial merger in the optical regime, followed by a rapid fade on the order of less than a week (Metzger et al., 2010), with longer wavelength emissions demonstrating a later peak and more gradual fade (Villar et al., 2017; Tanaka & Hotokezaka, 2013). In addition to the aforementioned electromagnetic (EM) signal, these mergers can also produce a $\gamma$-ray burst (GRB) (Tanvir et al., 2013; Berger et al., 2013; Fong & Berger, 2013; Yang et al., 2015) and are detectable via gravitational waves (GW) generated from the compact binary coalescence (CBC) (Rosswog, 2015; Nakar, 2020; Ascenzi et al., 2021). It was through detections by the International Gravitational Wave Network (IGWN) of the binary neutron star merger GW170817 (Abbott et al., 2017a) that the first (and only) KN associated with a GW signal, AT2017gfo, was localized and subsequently observed in a spate of multi-wavelength observations (Coulter et al., 2017; Smartt et al., 2017; Abbott et al., 2017a) in addition to the associated GRB170817A (Goldstein et al., 2017; Savchenko et al., 2017; Abbott et al., 2017b).
@@ 4 Data Processing
In the aggregate, all of the transient types in this study are characterized by a rise followed by a fall. Due to the nature of transient observation cadences, a smaller number of objects were observed during their rise compared to the number observed during their fade. Of the objects that passed the GP quality requirement, there are 3 objects observed only during their rise, 33 objects observed only during their fade, and 30 objects observed for both their rise and fade. The distribution of object types in the sample is shown in Figure 2. A brief description of each of these classes can be found in Appendix A.
@@
Observational data is cut to center around the rise, peak, and fade. Of the 33 objects with an observed rise, 8 have a rise duration that exceeds 10 days in any filter; of these, 5 are SNe IIb and 3 are SNe Ia. For the purposes of this paper, we measure the fade as the difference in magnitude between the peak luminosity and observations over the next 2 days. KNe, as fast fading objects, would generally only be visible for a few days following their peak. An object that is still observable by ZTF for weeks following its peak is unlikely to be a KN and can be excluded from consideration, but the longevity of a transient cannot always be ascertained from a handful of early-time observations.
@@ 4.1 Gaussian Process
The GP is applied to the first 50 days of data for all objects in our sample [^2] We restrict our analysis to the first 50 days because this study largely focuses on the rapid rise and fade of transients, making longer term data unnecessary.. This results in smooth light curves with high observing cadences, allowing for more precise analysis of rapid changes in luminosity. From there, each of the objects was reduced to only include observations at and around the peak luminosity, and the flux values from the GP are converted back to magnitude space by inverting Equations 1 and 2. An example of the GP applied to AT 2017gfo is shown in Figure 3.
After applying the GP to the 70 objects with classifications and redshifts, 4 of the transients did not have sufficient data to be modeled by the GP; an example of one such object can be seen in Figure 4, which the GP models as essentially flat. This is due to the relatively low number of observations, which take place over a single night. The third data point in the g-band suggests there may be an observed fade, but the error is too large to say this definitively. Regardless, this shows that there is a lower limit in terms of observations below which the GP cannot effectively interpolate the data.
Conversely, Figure 5 shows a IIb SN, ZTF20actqnhg, that was well fit by the GP. While this is more observed than ZTF20abwysqy, this is an unevenly sampled lightcurve with a somewhat ambiguous peak in the g-band and a paucity of observations immediately following the peak in the r-band. These properties would make calculating several metrics, including the color at peak and FWHM, less consistent based solely on the observed data points. The GP creates a plausible interpolation of the observed data, which enables more consistent calculations of metrics like the fade rate and FWHM.
As a result, the total number of objects included in the study is 66, with 11 different transient types represented in the sample. The distribution of types in the sample is shown in Figure 2. A full list of objects can be found in Table LABEL:tab:objects.
@@ 5 Evaluation Metrics
We compare the sample of transients to a grid of kilonova models computed with the 3D Monte Carlo radiative transfer code POSSIS (Bulla, 2019, 2023). POSSIS can handle arbitrary 3D geometry for the ejecta and simulate KN flux and polarization spectra for different observer viewing angles. Thermal radiation is represented by Monte Carlo packets that are assigned initial energies and frequencies and interact with matter while traveling through the expanding and opaque ejecta. In particular, we use the grid presented in (Dietrich et al., 2020b) [^3] The grid is publicly available at https://bit.ly/possis_models. The ejecta are assumed to be axially symmetric and are divided into two main components: a disk-wind ejecta component in the inner region (i.e., low velocities) and a dynamical ejecta component in the outer regions (i.e., high velocities) characterized by lanthanide-rich compositions close to the merger plane and lanthanide-poor compositions at polar angles. The masses of the two components, $m_{\rm wind}\in[0.01,0.03,0.05,0.07,0.09,0.11,0.13]\,M_{\odot}$ and $m_{\rm dyn}\in[0.001,0.005,0.01,0.02]\,M_{\odot}$, and the half-opening angle of the lanthanide-rich component, $\phi\in[0,15,30,45,60,75,90]^{\circ}$, are varied to create a grid of 196 models, with observables extracted for $11$ different viewing angles. All models share the same velocity structure, with the wind ejecta extending from 0.025 to 0.08c and the dynamical ejecta from 0.08 to 0.3c. The values for the model grid are generally plotted as black points in all figures.
We also calculate these metrics for AT2017gfo and include it under the “AT2017gfo” label. We note that, since ZTF was not operating at the time of its detection, there are no ZTF observations of AT2017gfo. Instead, we use observations from Pan-STARRS (Chambers et al., 2016), which captured the fade in 3 bands similar to those of ZTF [^4] We chose to use Pan-STARRS data for AT2017gfo due to the number of observations, from which the GP benefits. We elected to not include two other prominent KNe asssociated with GRBs 211211A (Rastinejad et al., 2022) and 230307A (Levan et al., 2023; Yang et al., 2024) in our analysis. In the former case, there are not enough Pan-STARRS observations to meaningfully fit with the GP, and in the latter case, it was not observed by Pan-STARRS.. The same GP is applied to these observations, which can be seen in Figure 3. Across all the metrics evaluated, AT2017gfo generally falls within the same range as the KN model grid.
In addition to the following sections, which consider a combination of metrics, we evaluate the numerical separation between transient classes and the model grid with respect to each metric using a ”distance” method that is described in Section 5.4.
@@ 5.1 Rise and Fade Rates
Figure 8: Fade rate vs/ peak luminosity in each filter for the transient sample compared to the KNe model grid. Generally more objects were observed during their fade as compared to their rise (See Figure 7). Most of the transients are clustered just above the KNe model grid with comparable fade rates. The model grid demonstrates two populations distinguished by fade rate, with a gap between the faster and slower fading populations that is most noticeable in the g-band. This bimodality is driven by a difference in the temperature parameter; the faster fading population is made up of simulated lightcurves with a lower temperature one day after the initial coalescence.
@@ 5.3 Full Width at Half Maximum
5.4 Metric Distance
$d=\frac{\sum_{i=0}^{n_{type}}\sum_{j=0}^{n_{grid}}|x_{i}-x_{j}|}{n_{type}\times n_{grid}}$ (3)
where $n_{type}$ is the number of objects contained within a given transient grouping, $n_{grid}$ is the number of simulated KNe in the model grid, $x_{i}$ is the metric value for a given transient, and $x_{j}$ is the metric value for a KNe in the model grid.
For visual clarity, we condense the imposters into 5 groups
  • GRB afterglows
  • IIb SNe
  • Misc SNe (Ia, Ib, Ibn, II, IIp)
  • Galactic transients (CVs, novae)
  • Other transients (FBOTs, TDEs)
Figure 12 shows radar plots for all of the metrics discussed in the preceding subsections. Take the first row, which corresponds to the peak luminosity, as a test case. The KN model has a fairly narrow range in terms of peak luminosity, so the self similarity value is almost zero. Across all filters, galactic sources show the highest value, which is logical given that these transients are known to have much lower peak absolute luminosities in comparison to KNe and other extragalactic transients; however, they can be a contaminant due to their relative proximity resulting in a high apparent magnitude. One should also note that, while it is fairly small, AT2017gfo has a value slightly above that of the KN model baseline, which is most pronounced in the g-band.
5.4.1 Rise and Fade Rates
5.4.2 Full Width at Half Maximum
The fourth row shows the FWHM for each filter. As shown in Figures 10 and 11, the range of FWHM values is fairly narrow for the model grid; correspondingly, the baseline metric distance value is relatively low. The transient classes demonstrating the greatest distance from the baseline are both categories of SNe, with IIb SNe being edged out by the other collected SNe category. SNe occur on a longer timescale in comparison to KNe, and this is reflected in Figure 10. The galactic and miscellaneous sources are also slightly above the baseline value, but much less so.
5.4.3 Color
@@ 6 Conclusions
While there is variance in the qualitative performance of each metric, most of the metrics discussed in this paper show a mild discriminatory utility; using these metrics in tandem with some sort of weighting could be used to calculate a second order metric, along the lines of the metric distance approach discussed in Section 5.4, to indicate the likelihood of a transient being a KN candidate. This would return results on a timescale shorter than inference-based methods that use tools like NMMA (Barna et al., 2024).
@@ A.1 GRB Afterglows
GRB afterglows (referred to as afterglows for much of this paper) occur as a result of the shock originating from a GRB interacting with the interstellar medium in the form of synchrotron emission (Sari et al., 1998). Like KNe, afterglows decline rapidly (Rhoads, 1999).
CBCs give rise to both a KN and a short GRB (sGRB) afterglow, as was observed with GW170817 and its counterparts (Abbott et al., 2017b). In these cases, it can be difficult to disentangle the two signals from each other given the greater brightness of afterglows (Zhu et al., 2022), though multi-wavelength observations can mitigate this challenge (Wallace & Sarin, 2024).
@@ A.5 Tidal Disruption Events
Tidal disruption events (TDEs) occur when a star orbiting a supermassive black hole (SMBH) is drawn so close to the SMBH that it is pulled apart by tidal forces. This phenomenon was first theorized in relation to quasi-stellar sources (QSOs) (Hills, 1975). While the first observations of TDEs occurred in the x-ray regime (Piro et al., 1988; Donley et al., 2002), TDEs are observable across the EM spectrum depending on the viewing angle (Dai et al., 2018). A number of TDEs have been detected by ZTF (van Velzen et al., 2021), making them a potential contaminant, but their presence could be mitigated with galaxy correlation given their presence at the center of galactic nuclei and their higher peak luminosity in comparison to KNe (Lodato & Rossi, 2011).
Some TDEs may however be found to be offset from their host galaxy nucleus. Three such cases have been reported in the literature (Lin et al., 2018; Jin et al., 2025; Yao et al., 2025), two of which were observed primarily in the X-rays (Lin et al., 2018; Jin et al., 2025). In the case of the optically-discovered off-nuclear event AT2024tvd (Yao et al., 2025), the transient reached maximum light ($M_{r}~{}-19$ mag) in $\sim 20$ days in the rest frame, then faded slowly. Its peak brightness and light curve are therefore distinct from any plausible kilonova models. Because of their rarity and light curve behavior, we expect contamination from off-nuclear TDEs to be negligible.
A.6 Type Ia Supernovae
There are two primary methods of delineating the various classes of SN, the first being by the presence of absorption lines of specific elements in their spectra and the second being their explosion mechanism. The former was initially described in Minkowski (1941) and has been adopted as the primary classification method and has been expanded in greater detail (da Silva, 1993). Broadly, Type I SNe are defined by the presence of a hydrogen absorption line, whereas Type II SNe do demonstrate a hydrogen absorption line.
Type Ia supernovae (SNe Ia) are differentiated by the presence of a silicon line in their spectra (Filippenko, 1997). They are also the only class of SN that undergoes thermonuclear runaway for its explosion mechanism. This occurs when a WD in a binary system accretes enough mass to hit the Chandrasekhar limit, which leads to the thermonuclear runaway (Chandrasekhar, 1931; Baym et al., 1971). Due to this roughly similar explosion mass, SNe Ia are understood to have standardizable peak luminosities (Phillips, 1993), though there are a subset of SNe Ia that are under-luminous, which are referred to as SNe Iax (Foley et al., 2013).
@@ A.6 Type Ia Supernovae
A.7 Type Ib Supernovae
Like SNe Ia, SNe Ib are categorized as Type I SNe due to the paucity of hydrogen in their spectra, but they demonstrate a helium line that is not present in SNe Ia and generally lack a strong silicon absorption line (Filippenko, 1997). SNe Ib are driven by core collapse (CC), which occurs when the core of the star is no longer able to counter the gravitational forces of its own mass (Clocchiatti & Wheeler, 1997). CC SNe progenitors are massive stars, generally of the order 8 $M_{\odot}$ at minimum (Heger et al., 2003; Díaz-Rodríguez et al., 2021), with some debate as to the upper mass limit (Smartt, 2009, 2015; Sukhbold et al., 2016; Kuncarayakti et al., 2018; Díaz-Rodríguez et al., 2021).
@@ A.8 Type Ibn Supernovae
A.9 Type II Supernovae
Like SNe Ib, SNe II occur as a result of the CC of massive stars, but they differ in that their red supergiant progenitors retain their hydrogen envelope and demonstrate a hydrogen absorption line in their spectra (Chevalier & Fransson, 1994; Filippenko, 1997). There are a number of subclasses of SNe II, of which two are described below.
@@ A.10 Type IIP Supernovae
A.11 Type IIb Supernovae
SNe IIb are a unique class of CC SNe that initially demonstrate spectra that appear similar to SNe II but later evolve to more closely resemble SNe Ib (Filippenko, 1997). This occurs due to a milder stripping of the outer hydrogen layers at early times in comparison to other SE SNe, with the associated hydrogen spectral features disappearing at later times, coming to resemble the spectra of SNe Ib (Sravan et al., 2019). A unique observational trait of SNe IIb is their double-peak feature (Crawford et al., 2025). The initial peak is powered by shock-cooling emission (Kamble et al., 2016; Armstrong et al., 2021), which later gives way to the secondary peak that is driven by the decay of 56Ni produced during the initial explosion (Arnett, 1980; Pellegrino et al., 2023). The first peak demonstrates a more rapid rise and fade than the second, but both peaks fade faster than something like a SN IIP. SNe IIb represent the largest subclass of observed SE SNe (Li et al., 2011a; Shivvers et al., 2017), though they are generally overrepresented in our sample relative to SNe IIP, which are the most observed subclass of SNe generally (Shivvers et al., 2017), suggesting they are a major contaminant in KNe searches.
Appendix B Transient List